State Bonnet’s mean value theorem for integrals. Apply it to show that:

|35 cosxdx/x|≤ 2/3


1
Expert's answer
2021-06-22T02:58:39-0400

BONNET MEAN VALUE THEOREM


If both f:[a,b] → R and g:[a,b]→ R Integrable on [a,b] and f is non negative and monotonically decreasing on [a,b] then there exist a point n ∈ [a,b] such that

"\int"ab f(x) g(x) dx = f(a) "\int"na g(x) dx

To prove the above indefinite integral,

Let f(x) = 1/x and g(x) = cos x

Here both f(x) and g(x) are integrable on [3,5]

and f(x) is non negative and monotonic decreasing on [3,5] .

So bonnet theorom is applicable here .

Now f(a) = f(3) = 1/3

"\int"53 cosx/x dx = 1/3 "\int"n3 cos x dx

= 1/3 ( sin n - sin 3)

Maximum value of sine function is 1

So (1+1)/3 = 2/3

Hence proved